OpenStudy (anonymous):

Show that the points A(0,2) ; B(3 , 1 ) ; C(0,-2) are vertices of right triangle.

OpenStudy (anonymous):

How do we solve this ?

OpenStudy (anonymous):

Are you just plotting them on a graph, so they are forming a triangle?

OpenStudy (anonymous):

I don't know it is my assignment question.

OpenStudy (anonymous):

yeah, im pretty sure you just graph them. if thats the case you just do, over none up 2, and label that one 'a' then do the rest over whatever up whatever. its as easy as that :)

OpenStudy (anonymous):

Put the points on a graph and you will see the answer...

OpenStudy (anonymous):

can you please tell me, if this is the right graph?

OpenStudy (anonymous):

Looks like a right angle to me....

OpenStudy (anonymous):

you could also use the converse of the pythagorean theorem... that is find the distance between all those points and see if they satisfy a^2+b^2=c^2

OpenStudy (anonymous):

the second point is not right, but please tell me if the procedure is right?

OpenStudy (anonymous):

So, dplanc it means I need to use distance formula here?

OpenStudy (anonymous):

Yes, it's right.

OpenStudy (anonymous):

No, that's a lot of unnecessary work when you can just see that it is right angled.

OpenStudy (anonymous):

@azib , yes... but i think you should only do that if you don't want to graph. i think what you're doing is ok.

OpenStudy (anonymous):

or even better, less work, just find the slopes between all three points... if m1*m2 = -1, you'll know two slopes are perpendicular...

OpenStudy (anonymous):

or even better, just look at the coordinates.....

OpenStudy (anonymous):

thank you @estudier and @dpaInc .. I think I should try both methods in my assignment.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

hey...